3. A wastewater treatment plant adds $700,000 to the cost of a computer chip manufacturing facility interest rate is 4%, and the end-of-life scrap value is $40,000? (6 points) due to capital expenditure for wastewater treatment? (3 points) a) What is the annualized capital cost (U) of this plant if its lifetime is 15 years, the b) If the facility manufactures 1 million chips per year, what is the added cost per chip
Answer #1

ANSWER:

A) Annualized capital cost = cost added to the computer chip manufacturing facility(a/p,i,n) - scrap value(a/f,i,n)

n = 15 years and i = 4%

Annualized capital cost = 700,000(a/p,4%,15) - 40,000(a/f,4%,15)

Annualized capital cost = 700,000 * 0.0899 - 40,000 * 0.0499

Annualized capital cost = 62,930 - 1,996

Annualized capital cost = 60,934

B) Added cost per chip = Annualized capital cost / no of chips = 60,934 / 1,000,000 = 0.0609

so the cost added per chip is $0.0609